OpenStudy (anonymous):

WILL REWARD!!! What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? point (11,3); slope 4/11 a.y-3=4/11(x+11) b.y-3=4/11(x-11) c.y-3=-4/11(x-11) d.y-11=4/11(x-3)

OpenStudy (anonymous):

So the equation for point slope form is y-y1=m(x-x1) (All u have to do is plug in the given values) m is the slope and x,y are your two coordinates y-3=4/11(x-11), so b is your answer :)

OpenStudy (anonymous):

thank you so much !

OpenStudy (anonymous):

no prob!